OpenStudy (anonymous):

Find a ⋅ b. a = 7i - 4j, b = 4i + 3j

OpenStudy (aroyni):

Hi

OpenStudy (aroyni):

=7*4(i.i)-4*3(j.j)=28-12=16

OpenStudy (aroyni):

does this make sense to you? it's just that i.j=0 and j.i=0

OpenStudy (anonymous):

I figured it out a long time ago, thanks though :)